OpenStudy (anonymous):

The half-life of a certain first order reaction is 15 minutes. What fraction of the original reactant concentration will remain after 2.0 hours??

OpenStudy (anonymous):

I know the half life formula for a 1st order is t=ln(2)/k

OpenStudy (anonymous):

The half-life is 15 minutes, so at time t = 15, you have half of the original amount left. \[\frac{1}{2}=e^{15k}\] (This comes from the half-life formula, which I'm assuming you know: \(y=Ce^{kt}\)). Solving for k yields \[k=\frac{\ln\left(\frac{1}{2}\right)}{15}\approx -0.046\] So, now you have to find the amount left after time = 2.0 hours = 120 minutes: \[y=Ce^{120k}\] Of course, you don't know what y and C are, but that's fine, since you actually are looking for the fraction of the original amount that remains, which is given by \[\frac{y}{C}=e^{120k}\] Solve for the LHS.

OpenStudy (anonymous):

what formula is the second one??

OpenStudy (anonymous):

You could use \(1/2\) as the base instead of \(e\).

OpenStudy (anonymous):

@Dama28 The second one isn't a formula, it's just solving for \(k\).

OpenStudy (anonymous):

but where did that come from? Because I've only been taught the integrated rate laws, the half life formulas and that's it.

OpenStudy (anonymous):

\[ \frac{1}{2}=e^{15k} \\ \ln\frac{1}{2}=\ln (e^{15k}) \\ \ln\frac{1}{2} = 15k \]

OpenStudy (anonymous):

I mean the y=Ce^120k

OpenStudy (anonymous):

Basically, suppose the original concentration is \(C_0\) We know that the concentration after one half-life is: \[ C_0(1/2) \]After two half-lives it is \[ C_0(1/2)(1/2) = C_0(1/2)^2 \]After \(n\) half-lives \[ C_0(1/2)^n \]Does this make sense so far?

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

We also know that one half-life takes 15 minutes, so to find the number of half-lives in minutes. So that means \(n = t/15\).

OpenStudy (anonymous):

So after \(t\) minutes, our concentration is: \[ C(t) = C_0(1/2)^{t/15} \]

OpenStudy (anonymous):

yes, I know how to get the 0.046, but I dont know what to do then

OpenStudy (anonymous):

After \(t\) minutes, the fraction of the remaining concentration is: \[ f(t) = \frac{C(t)}{C_0} =(1/2)^{t/15} \]

OpenStudy (anonymous):

but what would be the initial concentration?? that's my problem

OpenStudy (anonymous):

"What fraction of the original reactant concentration will remain after 2.0 hours??" Well this is just saying: If \(2.0 \times 60 = t\), what is \(f(t)\)?

OpenStudy (anonymous):

You don't need the initial concentration. They're asking for \(f(t)\) which fortunately doesn't contain the initial concentration.

OpenStudy (anonymous):

They don't want the actual concentration \(C(t)\) they want the fraction of concentration \(f(t)\)\[ f(t) = \frac{C(t)}{C_0} =\frac{C_0(1/2)^{t/15}}{C_0} = (1/2)^{t/15} \]

OpenStudy (anonymous):

I really dont understand, and Im pretty sure this is very simple.

OpenStudy (anonymous):

What don't you understand?

OpenStudy (anonymous):

Would I just multiply the k value I got (0.046) times 120 minutes??

OpenStudy (anonymous):

Are you expressing the equation as a base e or base 1/2?

OpenStudy (anonymous):

This is the formula I know and what I did: \[\ln \frac{ A _{t} }{ A _{0} }=-kt\]

OpenStudy (anonymous):

I plugged in the values of k=0.046 and t=120 minutes and got 5.54

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

now, should I find e^5.54??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I understand the problem now... you didn't want to read anything that is longer than 2 sentences.

OpenStudy (anonymous):

I get 255.9

OpenStudy (anonymous):

the answer is 1/256.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

You would have gotten that if you had just used \[ (1/2)^{t/15} \]Like I said.... http://www.wolframalpha.com/input/?i=%281%2F2%29%5E%28120%2F15%29

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

The problem is that you forgot to put a negative sign on exponent.

OpenStudy (anonymous):

are you kidding me? that's it?? wow!! I finally get it hahah

OpenStudy (anonymous):

Did you understand the formula I got though...?

OpenStudy (anonymous):

no, im sorry